OpenStudy (anonymous):

I have sone homework and i have some questions that i need help on. Is anybody willing to help

OpenStudy (anonymous):

@theopestudyowl

OpenStudy (amistre64):

it helps if youpost a question or two

OpenStudy (amistre64):

and define your difficulty with it

OpenStudy (anonymous):

Here u go

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

ok, and what are you having an issue with?

OpenStudy (amistre64):

if f(x) = g(x) then it must be true for all values of x we can do a really basic test to see if the options are equal to the given setup let x=0 .... and compare the results

OpenStudy (anonymous):

Can u help me understand it

OpenStudy (amistre64):

let x=0, and compare the results ... thats how you understand it

OpenStudy (amistre64):

not sure where the disconnect is ....

OpenStudy (amistre64):

if 2 things are equal, they are the same at every point. test out a point .. let x=0 (0)^3 + 8 = 8 --------------------- does: (0+8)^3 = 8 ? does: (0-2)(0+0+4) = 8 ? does: (0+2)(0+0+4) = 8 ?

OpenStudy (amistre64):

all this process does is eliminate some of the options. if see if the remaining option is truely equal, just distribute it thru and see if you can algebra it into x^3 + 8

OpenStudy (anonymous):

I did the work n m answer comes out t be no;no;yes

OpenStudy (anonymous):

@amistre64